the prime factors of 1058 are 2,2r-1,2r-1 Find "r"
Answer by checkley77(12844)

1058/2=529
sqrt529=23
2r-1=23
2r=23+1
2r=24
r=24/2
r=12 ans.
Proof:
2(2*12-1)(2*12-1)=1058
2(24-1)(24-1)=1058
2*23*23=1058
1058=1058